OpenStudy (anonymous):

5h – 9 = –16 + 6h 4 –7 7 10 4x + 4 = 9x – 36 –8 –7 8 –3 Which of the following equations has an infinite number of solutions? 3x – 3 = –4x 2y + 4 – y = 16 7x + 5 = 4x + 5 + 3x 6y – 2 = 2(y – 1)

OpenStudy (anonymous):

haha yu know how to do these. just as the others.

OpenStudy (anonymous):

Ok the first answer is 7.

OpenStudy (anonymous):

Second is 8.

OpenStudy (anonymous):

3rd im stuck

OpenStudy (anonymous):

help!!

OpenStudy (anonymous):

in the third problem, if you simplify the equation and come up with something that is ALWAYS true, then the equation has infinite solutions...

OpenStudy (anonymous):

Ok- which one is it? I have tried every one-I get stuck.

OpenStudy (anonymous):

here are the equations: 3x – 3 = –4x 2y + 4 – y = 16 7x + 5 = 4x + 5 + 3x 6y – 2 = 2(y – 1) that first one seems to have a one solution: 3x-3=-4x 7x=3 x=3/7 (one solution) did you do the second one?

OpenStudy (anonymous):

WHoops! sorry--got distracted lol. Yes I did do the second one.

OpenStudy (anonymous):

the second one only has 1 solution also... how 'bout the third one?

OpenStudy (anonymous):

third one: 7x + 5 = 4x + 5 + 3x simplify this...

OpenStudy (anonymous):

Yea-and I think it has no solution.

OpenStudy (anonymous):

Oh sorry--ok hang on.

OpenStudy (anonymous):

yes it does have a solution...

OpenStudy (anonymous):

?? OK I am TOTALLY missing something here!!

OpenStudy (anonymous):

try solving for x here: 7x + 5 = 4x + 5 + 3x

OpenStudy (anonymous):

just tell me what u end up with....

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

Nope! I got it! 7x+5 = 7x+5! Haha--my sister helped me! :)

OpenStudy (anonymous):

Write an inequality for the word sentence: k is less than zero. k > 0 k ≥ 0 k < 0 k ≤ 0 answer choice c right?

OpenStudy (anonymous):

notice that no matter what x is, the left side will ALWAYS equal the right... so in this case you have infinite solutions

OpenStudy (anonymous):

Yep! :) :)

OpenStudy (anonymous):

Write an inequality for the word sentence: k is less than zero. k > 0 k ≥ 0 k < 0 k ≤ 0 answer choice c right?

OpenStudy (anonymous):

ok good. i am closing ths Q now